# AdvRepair: Provable Repair of Adversarial Attack

### v1

@@ 3.3 Patch allocation
Once a set of patch modules has been allocated to an input, it works in the whole robustness region of this input. For an input $x^{\prime}\notin\bigcup_{i=1}^{n}B(x_{i},r)$, we allocate it with $\tau(x^{\prime})$, a set of patch modules. After that, all the inputs $x\in B(x^{\prime},r)$ share the same set of patch modules as $x^{\prime}$, i.e., $F(x)=\sum_{j\in\tau(x^{\prime})}P_{j}(x)+N(x)$, instead of allocating $x$ with $\tau(x)$. In this sense, our defense is constructed before adversarial attacks.
3.4 Repair in a feature space
To repair an extremely large DNN with high input dimensionality, the memory and computational overheads associated with the linear relaxations obtained from formal verification are notably significant and, at times, prohibitive. Also, faced with a huge DNN, the precision of abstraction-based verification methods like DeepPoly is severely harmed during a large number of layer-by-layer propagation. Consequently, the key to adopting AdvRepair to large DNNs is to reduce the input dimensionality and the size of the neural network in formal verification.
@@ 3.4 Repair in a feature space
The dimensionality of a feature space is usually far smaller than that of the input layer. If we start with a feature space in employing DeepPoly, it does not need to calculate the abstraction of a great deal of neurons in the convolutional layers, thus getting its efficiency and precision enhanced.
To conduct repair in a given feature space with AdvRepair, we first give an approximation of the feature space on every $B(x_{i},r)$. Here we do not require that this approximation should be a sound abstraction of the real semantics of the feature space on $B(x_{i},r)$, because calculating such an over-approximation with high precision is quite time-consuming. Instead, we simply sample in $B(x_{i},r)$ and obtain the tightest box region that contains the behaviors of the samples in the feature space as the approximation. This approximation is not sound yet, but it does not mean that the inputs in $B(x_{i},r)$ whose behaviors in the feature space are not within this approximation will not be repaired by the corresponding patch module $P_{i}$, because the indicator $\mathcal{I}_{\mathcal{C}}$ still remains the same. Since the approximation of the feature space is not sound, the repair does not have a provable guarantee. It is also worth mentioning that, even if two properties with different ground truth labels have their feature spaces overlapping, our repair still works. In this situation, although two contradicted properties are involved in the training process, but they must correspond to two different robustness regions $B(x_{i},r)$, and thus different patch modules $P_{i}$. The correspondence of the robustness regions $B(x_{i},r)$ and the patch $P_{i}$ is always preserved, so $P_{i}$ is always working for repairing on $B(x_{i},r)$ with the correct classification label $\ell_{0}$.
After we obtain an approximation $X_{(i)}$ for each $B(x_{i},r)$ in the feature space, we employ Alg. 1, where the initialization of $D[i]$ in Line 5 is $(N_{l,L}+P_{i},X_{(i)})$ instead. Since the repair in the feature space is not provable, the safety violation loss function $\mathcal{L}(\varphi)$ for $\varphi=(N_{l,L}+P_{i},X_{(i)})$ in Def. 3 is modified to be
@@
Since we are here considering the feature space in AdvRepair, it is natural to take it into consideration the cosine similarity and the distance of the behaviors of $x^{\prime}$ and $x_{i}$ in the feature space. If the ground truth label of $x^{\prime}$ and $x_{i}$ matches, a high cosine similarity and a small distance of the behaviors of $x^{\prime}$ and $x_{i}$ in the feature space strongly indicate that the patch module $P_{i}$ for $x_{i}$ is effective in fixing adversarial attacks in $B(x^{\prime},r)$. In this case, the patches with the same ground truth label as $x^{\prime}$ are still alternative for $x^{\prime}$, but we allocate the ten patches with the maximum value of
$\frac{\langle N_{0,l}(x^{\prime}),N_{0,l}(x_{i})\rangle}{\|N_{0,l}(x^{\prime})\|_{2}\cdot\|N_{0,l}(x_{i})\|_{2}}\cdot\|N_{0,l}(x^{\prime})-N_{0,l}(x_{i})\|_{2}.$
@@
Repair in a feature space is an effective way to make AdvRepair scale on large DNNs. Although we sacrifice provability, its repair performance and generalization are extremely ideal in practical scenarios, which we will see in our experimental evaluation.
4 Experimental Evaluation
In this section, we evaluate AdvRepair by answering the following research questions:
RQ1:What is the overall performance of AdvRepair in repairing local robustness and correcting safety property violations?
RQ3: Does AdvRepair have scalability to repair large networks?
RQ4: What impact does the number of adversarial samples needed for repair have on the effectiveness of the repair process?
4.1 Setup
All the experiments are conducted on a machine equipped with an AMD EPYC 7763 64-Core Processor, 256 GB of memory, and an NVIDIA GeForce RTX 3090 with 24 GB of GPU memory. Each experiment has a timeout set to 10 000 seconds.  [^1] AdvRepair is available at an anonymous link: https://drive.google.com/drive/folders/1NU170UZ0XmYKgg6diKZc-XMcqsUAVEib?usp=drive˙link
Dataset. We conduct evaluations of our approach on three datasets: MNIST [35], CIFAR-10 [34] and ACAS Xu [69, 29]. On MNIST, we train two fully connected neural networks and one convolutional neural network. On CIFAR-10, we train a VGG19 [54] and a ResNet18 [20]. We assess our approach across 35 ACAS Xu DNNs, which comprises six hidden layers, each housing 50 neurons. As documented in [32], these DNNs are expected to satisfy Property-2 but exhibit violations. Detailed model structures and hyperparameters are in Appendix 0.C. For the local robustness repair task, we subject the DNNs trained on MNIST and CIFAR-10 to PGD [43], generating 50, 100, 200, 500 and 1 000 adversarial samples. For MNIST, the radius $r$ is set to be 0.05, 0.1 and 0.3, while for CIFAR-10 it is set to be $\frac{4}{255}$ and $\frac{8}{255}$. On ACAS Xu, our objective is to repair the violation of Property-2 while preserving the original performance of the ACAS Xu DNNs. We sample 500 counterexamples as the buggy inputs for repair.
Metrics. To assess generalization, we sample ten adversarial examples on each $B(x_{i},r)$ different from $x_{i}^{*}$ to form a generalization set $D_{\mathrm{g}}$ for MNIST and CIFAR-10, and use an independent set of 5 000 counterexamples for ACAS Xu. Besides these, we have an independent test set $D_{\mathrm{t}}$ of size 10 000, 10 000 and 5 000 for MNIST, CIFAR-10 and ACAS Xu, respectively. We employ AutoAttack [14] to attack the repaired DNN $F$ on $\{x_{i}\}_{i=1}^{n}$ and $D_{\mathrm{t}}$ with the same radius $r$. The metrics we use include repair success rate ($\mathrm{RSR}$), repair generalization rate ($\mathrm{RGR}$), drawdown ($\mathrm{DD}$), defense success rate ($\mathrm{DSR}$), and defense generalization success rate ($\mathrm{DGSR}$), defined as follows:
$\mathrm{RSR}$ =|$=\frac{|\{\,i\mid C_{F}(x_{i}^{*})=C_{N}(x_{i})\,\}|}{n},\qquad\mathrm{RGR}=\frac{|\{{\,x\in D_{\mathrm{g}}}\mid C_{F}(x)=\ell_{x}\,\}|}{\left|D_{\mathrm{g}}\right|},$
$\mathrm{DSR}$ =|$=\frac{|\{\,i\mid\exists x\in\mathrm{AA}(F,B(x_{i},r)),C_{F}(x)\neq C_{N}(x_{i})\,\}|}{n},$
$\mathrm{DGSR}$ =|$=\frac{|\{\,x\in D_{\mathrm{t}}\mid\exists x^{\prime}\in\mathrm{AA}(F,B(x,r)),C_{F}(x^{\prime})\neq C_{N}(x)\,\}|}{\left|D_{\mathrm{t}}\right|},$
$\mathrm{FDD}=\frac{|\{\,{x\in D_{\mathrm{t}}}\mid C_{F}(x)=C_{N}(x)\,\}|}{|D_{\mathrm{t}}|}.$
Baselines. We compare AdvRepair with the state-of-the-art repair methods including CARE [61], APRNN [63], PRDNN [60], REASSURE [22] and ART [39]. Since APRNN and PRDNN require selecting a layer to repair on, we traverse all eligible layers across each baseline, and report the layer that exhibit the best performance. Additionally, for the local robustness repair task, we compare our approach with an adversarial training method TRADES [83], where we select the trained model with the best performance in 200 epochs with their default parameters.
Table 1: Results of repairing local robustness, where REASS and TRADE denote REASSURE and TRADES, respectively, and – means timeout or memory overflow.
@@ 4.1 Setup
CARE PRDNN APRNN REASS TRADE ART Ours CARE PRDNN APRNN REASS TRADE ART Ours CARE PRDNN APRNN REASS TRADE ART Ours
@@
4.2 Repair performance
For the overall repair performance, we focus on $\mathrm{RSR}$, $\mathrm{DD}$ and running time, which indicate whether the buggy behaviors of $\{x_{i}^{*}\}_{i=1}^{n}$ can be fixed, whether the accuracy is maintained, and the efficiency, respectively.
The results of the local robustness repair task on MNIST are summarized in Table 1. In the experiments of repairing local robustness, PRDNN, REASSURE and AdvRepair all achieve a perfect 100% $\mathrm{RSR}$, attributed to their provable constructions. APRNN, while provable, exhibits limitations in effectively repairing specific instances within the FNN_small model. Despite being a provable method, ART is less effective in repairing tasks involving high-dimensional inputs. CARE and TRADES, which do not have provable guarantee, demonstrate lower $\mathrm{RSR}$ compared to the provable approaches. CARE exhibits suboptimal repair performance for all the three models, possibly due to the intricate relationship between local robustness repair and the numerous parameters in the DNNs. The challenge lies in resolving such errors through fault localization and modifying part of parameters. While TRADES exhibits better repair results on the convolutional neural network compared to the fully connected ones, its efficacy diminishes when addressing local robustness errors with a larger radius. Concerning $\mathrm{DD}$, AdvRepair and REASSURE outperform the other methods on all the models.
Table 2: Results of correcting violation of safety properties on ACAS Xu
@@ 4.2 Repair performance
CARE PRDNN REASS ART Ours CARE PRDNN REASS ART Ours CARE PRDNN REASS ART Ours CARE PRDNN REASS ART Ours
@@
The results of correcting safety property violations on ACAS Xu are presented in Table 2. For the task of correcting safety property violation, AdvRepair, PRDNN, REASSURE, and ART successfully construct a provable repair, resulting in a 100% $\mathrm{RSR}$. In contrast, CARE attains an average $\mathrm{RSR}$ of 95.8%. Regarding $\mathrm{RGR}$, all baseline methods exhibit notable performance, with ART and our approach showcasing the highest performance at 100%. In terms of $\mathrm{FDD}$, our approach achieves the most favorable performance with zero, surpassing CARE, ART, PRDNN, and REASSURE. The results highlight the effectiveness of our approach in achieving provable repairs and maintaining high fidelity.
Answer to RQ1: AdvRepair consistently outperforms the baselines in both local robustness repair and safety property violation correction tasks. It achieves a high repair success rate coupled with lower drawdown, showcasing superiority over the alternative approaches. The efficiency of AdvRepair is acceptable.
4.3 Generalization
In this experiment, we evaluate the generalization of the repaired network and its defense capabilities against new adversarial attacks, as shown in Table 3. AdvRepair outperforms the other baselines in terms of $\mathrm{RGR}$, showcasing superior performance. Notably, PRDNN, APRNN and TRADES also demonstrate some level of generalization, especially on the convolutional neural network, while the remaining tools exhibit relatively poor results. In terms of $\mathrm{DSR}$, AdvRepair exhibits superior performance compared to the other tools. TRADES is the only alternative showing a certain defensive capability against adversarial attacks on FNN_big and CNN, particularly when the radius is small. When evaluating $\mathrm{DGSR}$, we find that AdvRepair once again achieves the best performance. CARE, PRDNN, APRNN and TRADES demonstrate limited defense capabilities against FNN_big and CNN, especially when the radius is small. Notice that $\mathrm{DGSR}$ is the most significant metric for assessing generalization to global inputs. These results are sufficient to show that AdvRepair achieves a breakthrough in repairing adversarial attacks.
Answer to RQ2: AdvRepair exhibits robustness defense against new adversarial attacks, surpassing the defensive performance of other tools. AdvRepair exhibits significant advantages in terms of the generalization of repairing adversarial attacks.
Table 3: Results of generalization and defense against new adversarial attacks
@@ 4.3 Generalization
CARE PRDNN APRNN REASS TRADE ART Ours CARE PRDNN APRNN REASS TRADE ART Ours CARE PRDNN APRNN REASS TRADE ART Ours
FNN_ small 0.05 50 7.6 44.4 47.2 10.8 80.0 10.0 100.0 0.0 0.0 2.0 0.0 26.0 6.0 100.0 8.9 3.2 2.3 0.0 9.4 8.9 90.0
100 2.2 48.7 46.7 9.1 0.0 10.0 100.0 1.0 0.0 0.0 0.0 2.0 20.0 100.0 9.3 5.1 0.8 0.0 9.8 11.3 96.4
@@
500 1.3 79.6 73.3 12.6 100.0 0.0 0.0 0.0 0.0 100.0 0.0 0.0 0.0 98.3
1000 1.1 85.9 70.6 16.0 100.0 0.0 0.0 0.0 0.0 100.0 0.0 0.0 0.0 98.2
4.4 Scalability
We assess the scalability of AdvRepair on VGG19 and ResNet18 for local robustness repair using CIFAR-10 through the feature-space repair. The experimental results are presented in Table 4. Regarding $\mathrm{RSR}$, AdvRepair, PRDNN and APRNN achieve 100%. TRADES also performs commendably, although it slightly trails behind the provable tools. Notably, CARE exhibits an inability to repair the local robustness properties on the CIFAR-10 dataset. In terms of $\mathrm{RGR}$, AdvRepair demonstrates superior performance, with TRADES also achieving better results than other tools. As for $\mathrm{DD}$, AdvRepair outperforms the others except in scenarios where the number of buggy input is 1 000. In this situation, TRADES attains the best performance. This phenomenon is reasonable, since TRADES tends to perform better in the DNN accuracy when the number of samples in adversarial training are getting large. Looking at $\mathrm{DSR}$, we find that AdvRepair exhibits significantly better performance than the other tools, with the latter lacking the capability to defend against adversarial attacks on the original samples. For $\mathrm{DGSR}$, AdvRepair achieves the same level as the accuracy, surpassing the other tools. Although TRADES also performs comparatively well, it remains notably lower than AdvRepair. The running time of AdvRepair is acceptable, regarding the far superior repair performance and generalization.
Answer to RQ3: AdvRepair demonstrates good scalability in repairing local robustness on large-scale DNNs, outperforming the other tools.
Table 4: Results of repairing local robustness on CIFAR-10

### v2

@@ III-C Patch allocation
III-D Repair in a feature space
Repairing large DNNs with high-dimensional inputs poses significant challenges due to the substantial memory and computational costs associated with formal verification techniques. These costs can often become prohibitive, especially when dealing with deep networks where the precision of abstraction-based methods, such as DeepPoly, deteriorates significantly over multiple layers of propagation. Consequently, the key to adopting PatchPro to large DNNs is to reduce the input dimensionality and the size of the neural network in formal verification.
@@ III-D Repair in a feature space
The dimensionality of a feature space is usually far smaller than that of the input layer. If we start with a feature space in employing DeepPoly, it does not need to calculate the abstraction of a great deal of convolutional layers, thus getting its efficiency and precision enhanced. Selecting the input feature space for repairing remains an open challenge, with limited theoretical research available. Our choice of the second-to-last fully connected layer is heuristic, based on the common understanding that earlier layers in a neural network are primarily responsible for feature extraction, while later layers handle decision-making. Repairing in the decision layer is more efficient and incurs lower abstraction costs.
To conduct repair in a given feature space with PatchPro, we first give an approximation of the feature space on every $B(x_{i},r)$. Here we do not require that this approximation should be a sound abstraction of the real semantics of the feature space on $B(x_{i},r)$, because calculating such an over-approximation with high precision is quite time-consuming. Instead, we simply sample in $B(x_{i},r)$ and obtain the tightest box region that contains the buggy behaviors of the samples in the feature space as the approximation. To identify samples in the feature space that are “close” to adversarial examples, we use the Projected Gradient Descent (PGD) method and Fast Gradient Sign Method (FGSM) to find samples in $B(x_{i},r)$, and obtain their corresponding feature space behaviors. Formally, the abstraction $(X_{(i)})_{j}$ can be defined as follows.
Definition 6.
Let $S$ be the set of samples obtained by applying PGD and FGSM to $N$ in $B(x_{i},r)$. For each sample $s\in S$, we compute its feature space behavior $N_{0,l}(s)$ at the $l$-th layer. The abstraction of $S$ in the feature space is defined as:
$(X_{(i)})_{j}=[\min_{s\in S}(N_{0,l}(s))_{j},\max_{s\in S}(N_{0,l}(s))_{j}]$
where $(X_{(i)})_{j}$ is the range of $j$-th dimension of $X_{(i)}$, $(N_{0,l}(s))_{j}$ is the $j$-th element of the $N_{0,l}(s)$, and $j=1,\ldots,n_{l}$.
These samples, being actual adversarial examples in real-world scenarios, provide direct guidance for feature space repair. This approximation is not sound yet, but it does not mean that the inputs in $B(x_{i},r)$ whose behaviors in the feature space are not within this approximation will not be repaired by the corresponding patch module $P_{i}$, because the indicator $\mathcal{I}_{\mathcal{C}}$ still remains the same. Since the approximation of the feature space is not sound, the repair does not have a provable guarantee. It is also worth mentioning that, even if two properties with different ground truth labels have their feature spaces overlapping, our repair still works. In this situation, although two contradicted properties are involved in the training process, but they must correspond to two different robustness regions $B(x_{i},r)$, and thus different patch modules $P_{i}$. The correspondence of the robustness regions $B(x_{i},r)$ and the patch $P_{i}$ is always preserved, so $P_{i}$ is always working for repairing on $B(x_{i},r)$ with the correct classification label $\ell_{0}$.
After obtaining an approximation $B_{S}$ for each $B(x_{i},r)$ in the feature space, we employ Alg. 1, where the initialization of $D[i]$ in Line 5 is $(N_{l,L}+P_{i},X_{(i)})$ instead. Since the repair in the feature space is not provable, the safety violation loss function $\mathcal{L}(\varphi)$ for $\varphi=(N_{l,L}+P_{i},X_{(i)})$ in Def. 4 is modified to be
@@
In performing feature layer repairs on large-scale networks, we persist in utilizing the $\tau$ function guided by the original network output to allocate corresponding repair modules. While the approximation $X_{(i)}$ is indeed effective in capturing commonality in buggy behaviors, establishing a direct association between the input neighborhood $B(x^{\prime},r)$ of a newly encountered point $x^{\prime}$ and its distribution within the feature layers of the extensive network proves challenging. Consequently, assigning repair modules based on the input neighborhood of $x^{\prime}$ emerges as a more targeted and reliable approach.
@@
Repair in a feature space is an effective way to make PatchPro scale on large DNNs. Although we sacrifice provability, its repair performance and generalization capabilities demonstrate remarkable effectiveness in practical scenarios, which we will see in our experimental evaluation.
TABLE I: Results of repairing local robustness. We represent each tool’s name with their first three letters, such as CAR for CARE, APR for APRNN, etc. And “–” means timeout or memory overflow.
@@
CAR PRD APR REA TRA ART Ours CAR PRD APR REA TRA ART Ours CAR PRD APR REA TRA ART Ours
@@
TABLE II: Results of correcting violation of safety properties on ACAS Xu
@@
CAR PRD REA ART Ours CAR PRD REA ART Ours CAR PRD REA ART Ours CAR PRD REA ART Ours
@@
IV Experimental Evaluation
In this section, we evaluate PatchPro by answering the following research questions:
RQ1: What is the overall performance of PatchPro in repairing local robustness and correcting safety property violations?
RQ3: Does PatchPro have scalability to repair large networks?
RQ4: In the context of PatchPro, how does the size and quantity of patch modules influence the efficacy of the repair?
IV-A Setup
All the experiments are conducted on a machine with an AMD EPYC 7763 64-Core Processor, 256 GB of memory, and an NVIDIA GeForce RTX 3090 with 24 GB of GPU memory. Each experiment has a timeout set to 10 000 seconds.
Dataset
We conduct evaluations on four common datasets: MNIST [29], CIFAR-10 [30], Tiny ImageNet[31], and ACAS Xu [32, 33]. The first three are widely recognized benchmarks in the field for studying neural network robustness [34, 4], and ACAS Xu is a commonly used benchmark in research on neural network verification and repair [35, 22, 20]. On MNIST, we use two fully connected networks and one convolutional network, while on CIFAR-10, we train a VGG19 [36] and a ResNet18 [37]. On Tiny ImageNet, we train a Resnet152 and a WRN101-2 [38]. We assess our approach on 35 ACAS Xu DNNs, which, as documented in [35], are expected to satisfy Property-2 but exhibit violations. For the local robustness repair task, we generate adversarial samples using PGD [4] for DNNs trained on MNIST and CIFAR-10 (50, 100, 200, 500, and 1,000 samples), as well as Tiny ImageNet (500 and 1,000 samples). The radius $r$ is set to 0.05, 0.1, and 0.3 for MNIST, $\frac{4}{255}$ and $\frac{8}{255}$ for CIFAR-10, and $\frac{2}{255}$ and $\frac{4}{255}$ for Tiny ImageNet. On ACAS Xu, we aim to repair the violation of Property-2 using one patch module while preserving the original performance. Although PatchPro repairs safety properties without requiring sample information, due to the needs of other tools such as CARE, we sample 500 counterexamples as the faulty inputs for repair. For CIFAR-10, the accuracies of VGG19 and ResNet18 are 93.4% and 88.3%, respectively. For Tiny ImageNet, the accuracies of WRN101-2 and ResNet152 are 64.4% and 68.2%, respectively. The maximum number of iterations $M$ in Alg. 1 is set to 25. In Alg. 2, the maximum number of epochs $R$, learning rate $\eta$, and selection number $K$ are set to 10, 10, and 800, respectively. All patch modules consist of a single linear layer, unless otherwise specified. On MNIST, the patch module takes the sample itself as input, while on CIFAR-10 and Tiny ImageNet, it takes the output from the network’s penultimate layer as input. For sampling the feature space, we use PGD attack with a step size of $\frac{2}{255}$ for 10 rounds, with each round consisting of 50 steps. We collect all adversarial examples generated during this process. Additionally, we use FGSM attack to generate 50 adversarial examples, which are then combined with the samples obtained from PGD. This combined set of samples is used for sampling the feature space.
Baselines
We compare PatchPro with the state-of-the-art repair methods including CARE [8], APRNN [22], PRDNN [20], REASSURE [21] and ART [25]. Since APRNN and PRDNN require selecting a layer to repair on, we traverse all eligible layers across each baseline, and report the layer that exhibit the best performance. Additionally, for the local robustness repair task, we compare our approach with an adversarial training method TRADES [39], where we select the trained model with the best performance in 200 epochs with their default parameters.
Metrics
To assess generalization, we sample 10 adversarial examples on each $B(x_{i},r)$ different from $x_{i}^{*}$ to form a generalization set $D_{\mathrm{g}}$ for MNIST, CIFAR-10 and Tiny ImageNet, and use an independent set of 5 000 counterexamples for ACAS Xu. Besides these, we have an independent test set $D_{\mathrm{t}}$ of size 10 000, 10 000, 10 000 and 5 000 for MNIST, CIFAR-10, Tiny ImageNet and ACAS Xu, respectively. We employ AutoAttack [5] to attack the repaired DNN $F$ on $\{x_{i}\}_{i=1}^{n}$ and $D_{\mathrm{t}}$ with the same radius $r$. The metrics we use include repair success rate ($\mathrm{RSR}$), repair generalization rate ($\mathrm{RGR}$), drawdown ($\mathrm{DD}$), defense success rate ($\mathrm{DSR}$), and defense generalization success rate ($\mathrm{DGSR}$), defined as:
$\mathrm{RSR}$ =|$=\frac{|\{\,i\mid C_{F}(x_{i}^{*})=C_{N}(x_{i})\,\}|}{n},$
$\mathrm{RGR}$ =|$=\frac{|\{{\,x\in D_{\mathrm{g}}}\mid C_{F}(x)=\ell_{x}\,\}|}{\left|D_{\mathrm{g}}\right|},$
$\mathrm{DSR}$ =|$=\frac{|\{\,i\mid\forall x\in\mathrm{AA}(F,B(x_{i},r)),C_{F}(x)=C_{N}(x_{i})\,\}|}{n},$
$\mathrm{DGSR}$ =|$=\frac{|\{\,x\in D_{\mathrm{t}}\mid\forall x^{\prime}\in\mathrm{AA}(F,B(x,r)),C_{F}(x^{\prime})=C_{N}(x)\,\}|}{\left|D_{\mathrm{t}}\right|},$
$\mathrm{FDD}=\frac{|\{\,{x\in D_{\mathrm{t}}}\mid C_{F}(x)\neq C_{N}(x)\,\}|}{|D_{\mathrm{t}}|}.$
TABLE III: Results of generalization and defense against new adversarial attacks
@@ IV-A Setup
CAR PRD APR REA TRA ART Ours CAR PRD APR REA TRA ART Ours CAR PRD APR REA TRA ART Ours
FNN_ small 0.05 50 7.6 44.4 47.2 10.8 80.0 10.0 100.0 0.0 0.0 2.0 0.0 26.0 6.0 100.0 8.9 3.2 2.3 0.0 9.4 8.9 91.3
100 2.2 48.7 46.7 9.1 0.0 10.0 100.0 1.0 0.0 0.0 0.0 2.0 20.0 100.0 9.3 5.1 0.8 0.0 9.8 11.3 96.6
@@
500 1.3 79.6 73.3 12.6 100.0 0.0 0.0 0.0 0.0 100.0 0.0 0.0 0.0 0.0 98.3
1000 1.1 85.9 70.6 16.0 100.0 0.0 0.0 0.0 0.0 100.0 0.0 0.0 0.0 0.0 98.2
IV-B Repair performance
The results of the local robustness repair task on MNIST are summarized in Table I. We evaluate overall repair performance with $\mathrm{RSR}$, $\mathrm{DD}$, and runtime, representing fix success, accuracy retention, and speed, respectively. PRDNN, REASSURE, and PatchPro reach 100% $\mathrm{RSR}$ due to their provable designs. APRNN, despite being provable, underperforms on some FNN_small cases, while ART struggles with high-dimensional inputs. CARE and TRADES, lacking provable guarantees, show inferior $\mathrm{RSR}$ compared to provable methods. CARE’s suboptimal repair across all models may stem from the complex interplay between local robustness repair and DNNs’ numerous parameters, making it difficult to pinpoint and modify relevant parts. Although TRADES performs better on CNNs, its effectiveness wanes with larger-radius local robustness errors. In terms of $\mathrm{DD}$, PatchPro and REASSURE consistently outperform alternative methods across all models.
The results of correcting safety property violations on ACAS Xu are presented in Table II. Each patch module here is of the same size as the original network. PatchPro, PRDNN, REASSURE, and ART achieve a perfect 100% $\mathrm{RSR}$ with provable repairs. Both ART and our method showcase remarkable $\mathrm{RGR}$ scores, achieving a perfect 100% among all the evaluated tools. PatchPro also excels with a $\mathrm{FDD}$ of zero, outperforming CARE, ART, PRDNN, and REASSURE. These results emphasize our method’s strength in delivering provable repairs and preserving high-fidelity functionality.
TABLE IV: Results of repairing local robustness on CIFAR-10